3a-5b=1
2a+3b=7
With working out

Answer by mathbird(71)

3a-5b=1-----------eq1
2a+3b=7 ----------eq2
multipling the eq 1 with 3 and eq 2 with 5
9a-15b=3
10a+15b=35
adding the above two equations we get
19a=38
a=38/19
a=2
put the value in eq 1
3a-5b=1
3*2-5b=1
6-5b=1
-5b=1-6
-5b=-5
b=1